OpenStudy (anonymous):

What is the value of b2 - 4ac for the following equation? x 2 + 5x + 4 = 0 a) 9 b) 16 c) 25

OpenStudy (anonymous):

a = 1, b = 5, c = 4.

OpenStudy (anonymous):

x 2 + 5x + 4 = 0 A B C 1 5 4 so 5^2-4*1*4 25-16=?

OpenStudy (anonymous):

9 lol

OpenStudy (anonymous):

just remember a quadratic equation in standard for is ax^2+bx+c notice how the x's exponent decreases ax^2+bx^0+cx^0 x^0 is actually 1 but not when x=0 it's undefined

OpenStudy (anonymous):

wait where is A, B, and C coming from?

OpenStudy (anonymous):

it just the standard form

OpenStudy (anonymous):

just look at it like templates

OpenStudy (anonymous):

ohhhhh wow I feel stupid sry thanks

OpenStudy (anonymous):

like if you remember y=mx+b you can say ax+b=y y=ax+b but just tradionally y=mx+b and traditionally and more logically y=ax^2+bx+c